PHOTODISSOCIATION OF VIBRATIONALLY EXCITED CH_{3}Cl: MODIFICATION OF THE DISSOCIA-TION DYNAMICS

\underline{H. MARK LAMBERT} and PAUL J. DAGDIGIAN, Department of Chemistry, The Johns Hopkins University, Baltimore, MD 21218.

Vibrationally excited CH_{3}Cl is prepared by laser excitation in the fourth overtone C-H stretch band near 725 nm and photodissociated at fixed wavelengths near 240 nm which also probe either the H, Cl(^{2}P_{3/2}), or Cl(^{2}P_{1/2}) photofragments via (2+1) resonance enhanced multiphoton ionization (REMPI) in a time-of-flight mass spectrometer. Jet cooling of the CH_{3}Cl reveals two partially resolved bands in the overtone excitation spectrum which are assigned to [5,0,0] and [4,0,0] + 2ν_{2}. The product yields from the photolysis of vibrationally excited CH_{3}Cl show decided enhancement compared to photolysis of the ground state. Branching ratios of atomic products are also determined for photolysis of both excited and ground state CH_{3}Cl. The ratio of Cl(^{2}P_{1/2}) to Cl(^{2}P_{3/2}) is dramatically larger for photolysis of vibrationally excited vs. ground state CH_{3}Cl.